# ARKit LabelMaker: A New Scale for Indoor 3D Scene Understanding

### v1

@@ 4.4 Results
4.5 LabelMakerv2 is generalizing beyond ARKitScenes
In order to demonstrate the effectiveness of our LabelMakerv2, we utilize it to process two self-captured scenes—a kitchen and a fireplace—taken in a holiday cottage using an iPhone 12 Max. In Figure 2, we present the reconstructed colored scenes alongside their semantic segmentation. It demonstrates visual accuracy and plausibility, confirming the effectiveness of our pipeline.
Figure 2: Self-captured scenes and the semantic segmentation generated by LabelMakerv2. This figure shows visually appealing segmentation results, therefore illustrates the effectiveness of our pipeline. The color mapping is defined in original LabelMaker’s repository.
4.6 Limitations & Broader Impact
While we extend LabelMaker [33] with a better pointcloud pipeline, we leave out the part that generates 2D segmentation maps. The computational cost of the NeRF-based lifting over the whole ArKitScenes dataset is beyond our available resources. It would be an interesting future direction of research to explore if training 2D models on this data yields similar performance gains as it is the case for 3D models.
Furthermore, 20 scenes in ARKitScenes processing are excluded due to lack of pose data. Our LabelMakerv2 requires accurate poses, but future iterations could integrate techniques such as bundle adjustments to reconstruct missing pose data.
Like the original LabelMaker [33], also our improved pipeline does not have perfect accuracy. While [33] showed that the accuracy is on par with crowd-sourced human annotations, there is always a danger of introducing systematic mistakes when training on noisy labels. For safety critical applications, rigorous testing on accurately annotated data is even more important when using tools like ours to source training data.
Does large-scale pre-training with automatic labels show similar trends as it does for language and image generation tasks? The discussed results point in this direction, with a measurable improvement to different models when pretraining on ArKitLabelMaker. However, training on large-scale real-world data ‘only’ achieves test results on par with the current SOTA based on synthetic data of even larger scale. Our conclusion is that real-world data is much more effective than syntethic data, but even larger overall scale is necessary to push the performance beyong state-of-the-art. Our developed pipeline makes it easy to provide training data once more scans are available.
5 Conclusion
In this paper, we presented the largest, real-world 3D RGB-D dataset with dense semantic annotations. The dense annotations are automatically generated using an improved version of LabelMaker [33], which we dub LabelMakerv2. While these labels are automatically generated and therefore imperfect, we demonstrate their value to pre-training commonly used 3D semantic segmentation methods significantly improving the performance of existing models trained with traditional, self-supervised, or augmentation-heavy training strategies. This allows to draw parallels to recent advances in language and image generation, where scaling up the size of training data led to huge gains in performance. Therefore, we also provide an integration of a commonly available 3D scanning software for iOS to enable the usage of mobile devices to easily generate more data for training and evaluation.
Acknowledgement
We give our many thanks to Xiaoyang Wu, one of the authors of Point Transformer V3 [36], who gave us many advice on the training of PTv3.
[…]
Supplementary Material
A Dataset and code download availability
Except from labelmaker.org, our dataset is available on huggingface at https://huggingface.co/datasets/labelmaker/arkit_labelmaker with DOI of 10.57967/hf/2389. The pipeline code for LabelMakerV2 is available at https://github.com/cvg/LabelMaker/. The preprocessing and training code for MinkowskiNet model and PointTransformerV3 is documented in https://github.com/quantaji/labelmaker-mix3d and https://github.com/quantaji/LabelMaker-Pointcept.
B Head, common and tail split mIoU scores for ScanNet200
In Table B1, we present the statistics for head, common, and tail classes. Our ALS200-pretrained MinkowskiNet significantly outperforms in common class mIoU and the jointly trained version of PTv3 shows a large increase in tail class mIoU.
Method Training Data Validation Test
head common tail head common tail
MinkUNet [7]
vanilla ScanNet200 52.3 22.5 13.2 46.3 15.4 10.2
fine-tune (Ours) ALS200 → ScanNet200 53.9 24.2 12.5 49.0 19.4 9.4
co-training (Ours) ALS200 + ScanNet200 55.1 24.7 12.4 - - -
PTv3 [36]
vanilla ScanNet200 56.5 30.1 19.3 - - -
fine-tune (Ours) ALS200 → ScanNet200 58.6 33.0 23.8 - - -
fine-tune (Ours) ALC200 → ScanNet200 58.2 33.1 25.0 58.2 30.9 22.2
PPT [36] ScanNet200 + S3DIS + Structure3D → ScanNet200 - - - 59.2 33.0 21.6
PPT(Ours) ScanNet+ ScanNet200 + ScanNet++ + Structure3D + ALC 60.9 35.48 24.6 61.0 32.2 27.1
Table B1: ScanNet200 validation and test mIoU for head, common and tail classes. For MinkowskiNet, ARKit LabelMaker pre-trained network shows significant improvement on head and common classes. For PTv3, we see improvements across all three splits.
C Detailed process of applying LabelMaker to ARKitScenes
@@
Optimizing compute resource scheduling. ARKitScenes contains scenes of a wide range of sizes, spanning from a minimum of 65 frames to a maximum of 13796, and different parts of the pipeline scale differently with increasing scene size. To figure out the minimum resources requirements, we select 11 scenes of varied sizes uniformly distributed within the minimum and maximum range and record their resources usage. While most jobs are not sensitive to scene size and can suffice with a fixed resource allocation, the base models exhibit greater sensitivity to scene size. We interpolate resource needs with respect to scene size and summarize the empirical numbers into Table C2. Through this, we ensure that we request minimal-required resources, so that we have lowest job waiting time and less idle compute power.
@@
Table C2: Requested resources for each task. We report the average resources required by the individual steps of the LabelMakerv2 pipeline. The required cores, RAM, and GPU time varies across the different jobs. Through our job scheduling mechanism, we ensure that the required compute is optimially distributed across all jobs.
@@
C.1 Implementation Details

### v2

@@ B PTv3 Results on ScanNet++
C Tail Classes Performance of PTv3
We give a detailed plot of the number of correctly predicted tail class points on ScanNet200 validation set in Figure C2.
Figure C2: Correctly predicted tail class points on ScanNet200 validation set. We compare the number of correctly predicted points of tail class in ScanNet200 validation sets between PTv3 trained from scratch and the PTv3-PPT trained with our datasets. With our dataset, Point Transformer gains more ability to detect rase classes.
D Detailed process of applying LabelMaker to ARKitScenes
@@
Optimizing compute resource scheduling. ARKitScenes contains scenes of a wide range of sizes, spanning from a minimum of 65 frames to a maximum of 13796, and different parts of the pipeline scale differently with increasing scene size. To figure out the minimum resources requirements, we select 11 scenes of varied sizes uniformly distributed within the minimum and maximum range and record their resources usage. While most jobs are not sensitive to scene size and can suffice with a fixed resource allocation, the base models exhibit greater sensitivity to scene size. We interpolate resource needs with respect to scene size and summarize the empirical numbers into Table D2. Through this, we ensure that we request minimal-required resources, so that we have lowest job waiting time and less idle compute power.
@@
Table D2: Requested resources for each task. We report the average resources required by the individual steps of the LabelMakerv2 pipeline. The required cores, RAM, and GPU time varies across the different jobs. Through our job scheduling mechanism, we ensure that the required compute is optimially distributed across all jobs.
@@
E Log-Linear Performance Relation in Data Scaling
Figure E3: Relation of Validation mIoU against training data percentage of ARKit LabelMaker. This figure shows the validation mIoU on ScanNet200 after fine-tuning with respect to the percentage of ARKit LabelMaker data used in pre-training. This figure shows a log-linear relationship.
E.1 Implementation Details
We optimize the code to deploy each individual piece of the pipeline of LabelMakerV2 as individual jobs to a GPU cluster, with SLURM as a dependency manager between the pipeline pieces. To optimize the overall execution time, it is therefore important to be able to estimate the processing time of each piece of the pipeline at the point of job submission. ARKitScenes contains scenes of a wide range of sizes, spanning from a minimum of 65 frames to a maximum of 13796, and different parts of the pipeline scale differently with increasing scene size. To figure out the minimum resources requirements, we select 11 scenes of varied sizes uniformly distributed within the minimum and maximum range and record their resources usage. While most jobs are not sensitive to scene size and can suffice with a fixed resource allocation, the base models exhibit greater sensitivity to scene size. We interpolate resource needs with respect to scene size and summarize the empirical numbers in the Appendix. Through this, we ensure that we request minimal-required resources, so that we have lowest job waiting time and less idle compute power.